# Joint Time–Position Statistics and Fisher Information in Drift–Diffusion Molecular Channels

### v1

@@ Abstract
This letter presents a closed-form characterization of the joint distribution of first arrival time (FAT) and first arrival position (FAP) in diffusion-based molecular communication (MC) systems with drift. Prior studies have investigated FAT modeling via inverse Gaussian distributions [1] and applied FAT statistics for parameter estimation and synchronization tasks [2, 3], while more recent work has characterized FAP for spatial channel analysis [4]. In contrast, we derive an explicit joint probability density function (PDF) under constant drift and isotropic diffusion in arbitrary spatial dimensions. Our result reveals a nontrivial coupling between arrival time and lateral position, generalizing known inverse Gaussian models. We further compute the Fisher information matrix (FIM) with respect to key channel parameters, showing that the joint observation enables estimation of lateral drift and improves sensitivity to the diffusion coefficient—capabilities not achievable with time-only or position-only models. This joint framework enhances the modeling and inference capabilities for molecular communication channels where spatial randomness itself carries non-negligible information.
@@ I Introduction
We further analyze the Fisher information matrix (FIM) associated with the joint observation, quantifying its utility for parameter estimation. Prior work has applied FIM analysis to other MC settings [9, 2, 3], but not in conjunction with FAT–FAP statistics. Our results show that the joint model enables identification of lateral drift components and improves sensitivity to diffusivity—capabilities not achievable with FAT-only or FAP-only models.
This joint perspective provides a stronger foundation for inference and system design in MC applications, particularly in regimes where spatial randomness carries non-negligible information.
@@ II System Model
We consider a single-input single-output (SISO) molecular communication (MC) system in a $D$-dimensional unbounded fluid medium with constant drift and diffusion. The transmitter (TX) and receiver (RX) are modeled as parallel (D1)(D-1)-dimensional hyperplanes, separated by one unit length along the $x^{(1)}$ axis.
To simplify the analysis, we normalize length by the TX–RX distance $\lambda$ and time by $\lambda/v_{p}$, where $v_{p}>0$ denotes the perpendicular drift component. Under this normalization, the system becomes dimensionless; in particular, the perpendicular drift is scaled to $1$, so that the drift vector takes the form $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})$.
The TX emits a single molecule from location
@@
where $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})\in\mathbb{R}^{D}$ is the drift vector, $\sigma>0$ is the dimensionless diffusion coefficient, and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
@@
and the first arrival position as the random vector
@@
In the next section, we derive the joint distribution of $(T,\mathbf{X}_{T})$, which captures the spatiotemporal coupling of first-arrival dynamics under drift–diffusion.
@@ III Joint Time-Position Distribution of First Arrivals
We derive the joint probability density function (PDF) of the first arrival time $T$ and the first arrival position $\mathbf{X}_{T}$ of a diffusing particle under constant drift in a dimensionless setting. The transmitter is located on the hyperplane $\{x^{(1)}=0\}$, and the fully absorbing receiver lies on $\{x^{(1)}=1\}$. A molecule is released from
$\mathbf{x}_{0}=(0,x_{0}^{(2)},\ldots,x_{0}^{(D)})$ (5)
at time $t=0$, and its position $\mathbf{X}_{t}$ evolves according to the stochastic differential equation
where $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})\in\mathbb{R}^{D}$ is the drift vector, $\sigma>0$ is the diffusion coefficient, and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
The first arrival time is defined as
$T:=\inf\{t>0:X_{t}^{(1)}=1\},$ (7)
and the corresponding first arrival position is $\mathbf{X}_{T}=(1,X_{T}^{(2)},\ldots,X_{T}^{(D)})$. Throughout, we assume that the components of $\mathbf{B}_{t}$ are independent and that the receiver is absorbing only in the $x^{(1)}$-direction.
@@
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&\overset{(a)}{=}f_{T}(t)\,f_{\mathbf{X}_{T}|T}(\mathbf{x}|t)\\ &\overset{(b)}{=}f_{T}(t)\,f_{\mathbf{X}_{t}|T}(\mathbf{x}|t)\\ &\overset{(c)}{=}f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),\end{split}$ (8)
where (a) follows from the definition of the conditional density, (b) uses the equality in distribution $(\mathbf{X}_{T}|T=t)\stackrel{{\scriptstyle d}}{{=}}(\mathbf{X}_{t}|T=t)$, and (c) holds because the lateral process $\mathbf{X}^{(2:D)}_{t}:=(X_{t}^{(2)},\ldots,X_{t}^{(D)})$ is independent of $T$. This independence stems from the geometry of the problem: the hitting time $T$ is determined solely by the one-dimensional process $\{X_{t}^{(1)}\}_{t\geq 0}$, while the components of the Wiener process are mutually independent. This argument extends the proof of [10, Thm. 2.37] to arbitrary dimension $D$.
@@
where $f_{T}(t)$ is the inverse Gaussian (IG) density
@@
The marginal density of the first arrival position (FAP) is obtained by integrating over $t$:
@@
Identifying the resulting integral as a special case of the identity involving the modified Bessel function $K_{\nu}(\cdot)$ [11, Eq. (3.478.4)], we obtain
@@ IV Fisher Information Analysis
We next analyze the information contained in the joint observation $(T,\mathbf{X}_{T})$ with respect to key model parameters. In particular, we compute the Fisher information matrix (FIM) associated with the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, based on the joint PDF derived in Section III. The FIM characterizes the fundamental limits of parameter estimation accuracy for drift and diffusion in molecular communication systems.
From (12), the joint density can be written as
$f_{T,\mathbf{X}_{T}}(t,\mathbf{x})=\delta(x^{(1)}-1)\,(2\pi\sigma^{2})^{-D/2}\,t^{-D/2-1}\,\mathbbm{I}\{t>0\}$
$\quad\quad\times\exp\!\left(-\frac{(1-t)^{2}+\|\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\|^{2}}{2\sigma^{2}t}\right).$ (15)
The corresponding log-likelihood function (up to the support constraint $t>0$ and $x^{(1)}=1$) is
@@
For convenience, define the residual vector
@@ IV-B Fisher Information Matrix Entries
We first consider the entry corresponding to the diffusion coefficient $\sigma$. Differentiating the log-likelihood gives
@@
Taking expectations yields
$I_{\sigma\sigma}$ $=-\mathbb{E}\!\left[\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}T}\Big{(}(1-T)^{2}+\|\boldsymbol{\Delta}_{T}\|^{2}\Big{)}\right]$
@@
The two expectations can be computed separately. Since $T\sim\mathrm{IG}(1,1/\sigma^{2})$ with moments $\mathbb{E}[T]=1$ and $\mathbb{E}[T^{-1}]=1+\sigma^{2}$, it follows that
@@
Moreover, conditioned on $T=t$, the lateral displacement satisfies $\boldsymbol{\Delta}_{T}\sim\mathcal{N}(0,\sigma^{2}tI_{D-1})$, which gives
@@
Taking expectation over $T$ yields
$\mathbb{E}_{T}\!\left[\frac{1}{T}\,\mathbb{E}_{\mathbf{X}_{T}|T}[\|\boldsymbol{\Delta}_{T}\|^{2}|T]\right]=(D-1)\sigma^{2},$ (26)
@@
Taking expectations, we obtain
$I_{\sigma,v^{(k)}}=\frac{2}{\sigma^{3}}\,\mathbb{E}[\Delta_{T}^{(k)}]=0,$ (29)
showing that the FIM remains diagonal.
Finally, for each drift component $v^{(k)}$, the second derivative takes the form
@@
$I_{v^{(k)}v^{(k)}}=\frac{1}{\sigma^{2}}\,\mathbb{E}[T]=\frac{1}{\sigma^{2}}.$ (31)
@@
Figure 1: Empirical conditional joint PDF of $(T,X_{T}^{(2)})$ obtained from particle-based simulations.
Figure 2: Analytical conditional joint PDF of $(T,X_{T}^{(2)})$ computed from the closed-form expression.
IV-C Discussion
The closed-form expressions above show that the joint observation of first arrival time $T$ and position $\mathbf{X}_{T}$ provides valuable information on both drift and diffusion. Under pure vertical drift, the Fisher information $I_{v^{(k)}v^{(k)}}=1/\sigma^{2}$ is strictly positive, implying that lateral drift components $v^{(k)}$ are estimable from $\mathbf{X}_{T}$ but not from $T$ alone. Moreover, the diffusion term $I_{\sigma\sigma}$ reflects contributions from both temporal and spatial observations, indicating that the joint model carries richer information than either marginal model.
Thus, the joint spatiotemporal framework captures a more complete statistical structure than time-only or position-only observations. In practice, combining $T$ and $\mathbf{X}_{T}$ in estimation or detection can improve sensitivity to environmental parameters such as background flow direction or diffusivity heterogeneity.
@@ V Numerical Simulations
We perform numerical experiments to validate the theoretical channel model in (12). For illustration, we focus on the two-dimensional (2D) case with a linear absorbing receiver modeled as the infinite line $\{(x^{(1)},x^{(2)})\in\mathbb{R}^{2}:x^{(1)}=1\}$. In this setting, the arrival position always satisfies $X_{T}^{(1)}=1$, so it suffices to examine the joint distribution of $(T,X_{T}^{(2)})$.
[…]
Since the simulation time horizon is limited to $2$ seconds, the empirical joint PDF obtained should be interpreted as conditional, i.e., conditioned on arrival within the interval $[0,2]$ seconds. For a fair comparison, we therefore also compute the corresponding conditional version of the theoretical joint PDF from (12). The results, shown in Fig. 2 and Fig. 2, demonstrate excellent agreement between theory and simulation.
@@ VI Conclusion
We presented a closed-form expression for the joint distribution of the first arrival time and position of a molecule propagating in a drift–diffusion channel. This result generalizes the classical inverse Gaussian model by explicitly capturing the spatiotemporal coupling of molecular transport in arbitrary spatial dimensions.
Building on this joint distribution, we derived the Fisher information associated with key environmental parameters. The analysis shows that combining temporal and spatial observations enables estimation of lateral drift components and improves sensitivity to the diffusion coefficient—capabilities not attainable with time-only models.
Future work may relax the constant-drift assumption to more general non-uniform drift fields, where drift strength or direction varies spatially.

### v2

# Joint Time-Position Statistics and Fisher Information in Drift-Diffusion Molecular Channels

@@ Abstract
This letter derives a closed-form joint distribution of the first arrival time (FAT) and first arrival position (FAP) in diffusion-based molecular communication (MC) systems with drift. Unlike prior work that studied FAT or FAP separately, we obtain an explicit joint probability density function under constant drift and isotropic diffusion in arbitrary dimensions, revealing a nontrivial spatiotemporal coupling. Based on this result, we compute the Fisher information matrix (FIM) and show that joint observations enable estimation of lateral drift and improve sensitivity to diffusion—capabilities not attainable with time-only or position-only models.
@@ I Introduction
We further analyze the Fisher information matrix (FIM) associated with the joint observation, quantifying its utility for parameter estimation. Prior work has applied FIM analysis to other MC settings [9, 5, 6], but not in conjunction with FAT–FAP statistics. Our results show that the joint model enables identification of lateral drift components and improves sensitivity to diffusivity estimation—capabilities not achievable with FAT-only or FAP-only models.
This joint perspective provides a stronger foundation for parameter estimation and system design in MC applications, particularly in regimes where spatial randomness carries non-negligible information.
@@ II System Model
We consider a single-input single-output (SISO) molecular communication (MC) system in a $D$-dimensional unbounded fluid medium with constant drift and diffusion, where a transmitter (TX) releases molecules that are absorbed by a receiver (RX). We adopt a non-dimensionalized formulation: length is normalized by the transmitter–receiver (TX–RX) distance $\lambda$, and time is normalized by $\lambda/v_{p}$, where $v_{p}>0$ denotes the perpendicular [^1] Namely, $v_{p}$ is perpendicular to the parallel hyperplanes TX and RX. drift component. Under this scaling, the system becomes dimensionless; in particular, the perpendicular drift is scaled to $1$, so that the drift vector takes the form $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})$. Unless otherwise noted, all theoretical results in this section are expressed in these normalized units.
After normalization, the TX and RX are modeled as parallel (D1)(D-1)-dimensional hyperplanes separated by one unit length along the $x^{(1)}$ axis. The TX emits a single molecule from location
@@
where $\sigma>0$ is the dimensionless diffusion coefficient and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
@@
and the corresponding first arrival position as
@@ III Joint Time-Position Distribution of First Arrivals
We now derive the joint probability density function (PDF) of the first arrival time $T$ and first arrival position $\mathbf{X}_{T}$ under the drift–diffusion dynamics defined in Section II.
@@
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&=f_{T}(t)\,f_{\mathbf{X}_{T}|T}(\mathbf{x}|t)\\ &=f_{T}(t)\,f_{\mathbf{X}_{t}|T}(\mathbf{x}|t)\\ &=f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),\end{split}$ (5)
where the last equality holds because the lateral process $\mathbf{X}^{(2:D)}_{t}:=(X_{t}^{(2)},\ldots,X_{t}^{(D)})$ is independent of $T$. The hitting time is determined solely by $\{X_{t}^{(1)}\}$, while the Wiener components are mutually independent. This extends the argument of [10, Thm. 2.37] to arbitrary dimension $D$.
@@
where $f_{T}(t)$ is the $\text{IG}(1,1/\sigma^{2})$ density
@@
The marginal density of FAP is obtained as
@@
By applying the identity for the order-$\nu$ modified Bessel function $K_{\nu}(\cdot)$ [11, Eq. (3.478.4)], we obtain
@@ IV Fisher Information Analysis
We next analyze the information contained in the joint observation $(T,\mathbf{X}_{T})$ with respect to key model parameters. In particular, we compute the FIM associated with the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, based on the joint PDF derived in Section III. The FIM characterizes the fundamental limits of parameter estimation accuracy for drift and diffusion in molecular communication systems.
The log-likelihood function (up to the support constraint $t>0$ and $x^{(1)}=1$) corresponding to the joint PDF (9) is
@@
For convenience, define the residual vector at time $t$ as
@@ IV-B Fisher Information Matrix Entries
We first consider the entry corresponding to the diffusion coefficient $\sigma$. Differentiating (14) with respect to $\sigma$ gives
@@
From (17), we have
$I_{\sigma\sigma}:=[\mathbf{I}(\boldsymbol{\theta})]_{1,1}$ $=-\mathbb{E}\!\left[\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}T}\Big{(}(1-T)^{2}+\|\boldsymbol{\Delta}_{T}\|^{2}\Big{)}\right]$
@@
The two terms in the expectation can be computed separately. Since $T\sim\mathrm{IG}(1,1/\sigma^{2})$ with moments $\mathbb{E}[T]=1$ and $\mathbb{E}[T^{-1}]=1+\sigma^{2}$, it follows that
@@
Moreover, conditioned on $T=t$, the residual vector satisfies $\boldsymbol{\Delta}_{T}\sim\mathcal{N}(0,\sigma^{2}tI_{D-1})$, which gives
@@
Applying the law of total expectation then yields
$\mathbb{E}\!\left[\frac{\|\boldsymbol{\Delta}_{T}\|^{2}}{T}\right]=\mathbb{E}_{T}\!\left[\frac{1}{T}\,\mathbb{E}_{\mathbf{X}_{T}|T}[\|\boldsymbol{\Delta}_{T}\|^{2}|T]\right]=(D-1)\sigma^{2},$ (22)
@@
From (17), we have
$I_{\sigma,v^{(k)}}:=[\mathbf{I}(\boldsymbol{\theta})]_{1,k}=\frac{2}{\sigma^{3}}\,\mathbb{E}[\Delta_{T}^{(k)}]=0,$ (25)
and similarly, $I_{v^{(k)},\sigma}:=[\mathbf{I}(\boldsymbol{\theta})]_{k,1}=0$, showing that the FIM is diagonal.
Finally, for each drift component $v^{(k)}$, $k\in\{2,\ldots,D\}$, the second derivative takes the form
@@
$I_{v^{(k)}v^{(k)}}:=[\mathbf{I}(\boldsymbol{\theta})]_{k,k}=\frac{1}{\sigma^{2}}\,\mathbb{E}[T]=\frac{1}{\sigma^{2}}.$ (27)
@@
Figure 1: Empirical conditional joint distribution obtained from particle-based simulations. Here, the drift velocity vector is $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 2: Analytical conditional joint distribution computed from the closed-form expression. Here, the drift velocity vector is $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 3: Empirical conditional joint distribution obtained from particle-based simulations under drift velocity vector $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 4: Analytical conditional joint distribution derived from the closed-form expression under drift velocity vector $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
@@ V Numerical Simulations
In this section, we revert to dimensionalized physical parameters to align with common practice in molecular communication studies. All quantities are therefore reported in standard physical units (e.g., micrometers for distance, micrometers per second for velocity, and seconds for time).
We conduct numerical experiments to validate the theoretical channel model in (9). For illustration, we consider the two-dimensional (2D) case with a linear transmitter and a linear absorbing receiver in parallel, and examine the joint distribution of first-hitting time and position.
Particle-based simulations are performed by discretizing Brownian motion with drift via the Euler–Maruyama method [1, Eq. (9)–(11)], and the empirical joint PDF is estimated for comparison. The physical parameters are set as follows: TX–RX distance $\lambda=1\leavevmode\nobreak\ \mu\text{m}$, diffusion coefficient $0.5\leavevmode\nobreak\ \mu\text{m}^{2}/\text{s}$, and initial release position $(0,0)\leavevmode\nobreak\ \mu\text{m}$. Two drift velocity vectors are considered: $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ and $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$. Each simulation uses $10^{6}$ particles with $2\times 10^{3}$ time steps in the Euler–Maruyama scheme, where each step corresponds to $10^{-3}$ seconds.
Since the simulation horizon is limited to $2$ seconds, the empirical joint PDF should be interpreted as conditional, i.e., conditioned on arrivals within $[0,2]$ seconds. For fair comparison, we likewise compute the corresponding conditional theoretical joint PDF from (9). As shown in Fig. 22 for drift $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ (no lateral drift) and Fig. 44 for drift $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ (with lateral drift), the simulation results exhibit excellent agreement with the theoretical predictions.
@@ VI Conclusion
We derived a closed-form expression for the joint distribution of the first arrival time and position of a molecule propagating in a drift–diffusion channel. This result generalizes the classical inverse Gaussian model by explicitly capturing the spatiotemporal coupling of molecular transport in arbitrary spatial dimensions.
Our Fisher information analysis highlights that the joint spatiotemporal observation provides richer information than marginal time-only or position-only models. In particular, lateral drift components, which are not identifiable from first arrival time alone, become estimable when incorporating arrival position. Moreover, the diffusion term aggregates contributions from both temporal and spatial observations, demonstrating the statistical advantage of the joint model for parameter estimation tasks.
These findings suggest that combining temporal and spatial observations could enable more sensitive estimation of key environmental parameters such as background flow direction and diffusivity heterogeneity. Future work may extend this framework to non-uniform drift fields, where drift strength or direction varies spatially.
Acknowledgment
This work was partially supported by the National Science and Technology Council of Taiwan under Grant NSTC 113-2115-M-008-013-MY3. The authors also acknowledge the use of ChatGPT (OpenAI, GPT-5, Aug. 2025) for minor language editing. All technical content and conclusions are the sole responsibility of the authors.